OpenStudy (mathmath333):

A vessel is full of a mixture of kerosene and petrol in which there is 18% kerosene.Eight litres are drawn off and then the vessel is filled with petrol.If the kerosene is now 15% how much does the vessel hold?

OpenStudy (mathmath333):

a.40 b.36 c.48 d.32

OpenStudy (mathmath333):

@freckles

OpenStudy (mathmath333):

@amistre64

OpenStudy (amistre64):

how much what does it hold? total volume?

OpenStudy (mathmath333):

total litres it can have

OpenStudy (amistre64):

lets develope some inital conditions

OpenStudy (amistre64):

k + p = n k/(k+p) = .18 p/(k+p) = .82 subtract 8 and add in some L k/(k+p-8+L) = .15 p/(k+p-8+L) = .85

OpenStudy (mathmath333):

ok how to proceed now

OpenStudy (amistre64):

substitution comes to mind

OpenStudy (amistre64):

k/(k+p) = .18 k = .18k + .18p .82k =.18p k = 18/82 p which reduces this to equations of p and L

OpenStudy (mathmath333):

ok so how to proceed ,i cant get the way

OpenStudy (amistre64):

we know: k/(k+p-8+L) = .15 p/(k+p-8+L) = .85 and we know k = 9/41 p

OpenStudy (amistre64):

multiply acorss, and combine like terms, then you are set for a system of 2 equations in 2 unknowns that will hopefully give us a solution :)

OpenStudy (mathmath333):

m stii confused whick term to multiply with which term

OpenStudy (amistre64):

k/(k+p-8+L) = .15 p/(k+p-8+L) = .85 substitute for k ... show me some work

OpenStudy (mathmath333):

9/[41p(k+p-8+L)] = .15 p/(k+p-8+L) = .85 do i have to multiply this to cancel p

OpenStudy (amistre64):

k/(k+p-8+L) = .15 p/(k+p-8+L) = .85 k = .15(k+p-8+L) p = .85(k+p-8+L) k = .15k +.15p -8(.15) + .15L p = .85k + .85p -8(.85) +.85L 9/41 p = .15(9/41 p) +.15p -8(.15) + .15L p = .85(9/41 p) + .85p -8(.85) +.85L 0 = p(.15(9/41) + .15 - 9/41) -8(.15) + .15L 0 = .85(9/41 p) + .85p - p -8(.85) +.85L 0 = p(.15(9/41) + .15 - 9/41) -8(.15) + .15L 0 = p(.85(9/41) + .85 - 1) -8(.85) +.85L 8(.15) = p (.15(9/41) + .15 - 9/41) + .15L 8(.85) = p (.85(9/41) + .85 - 1) +.85L

OpenStudy (amistre64):

8(.15) = -3/82 p + .15L 8(.85) = 3/82 p + .85L hmmm it seems to reduce so that it eliminates p from the getgo

OpenStudy (amistre64):

im gonna have to make sure i kept things in order

OpenStudy (amistre64):

A vessel is full of a mixture of kerosene and petrol k+p = n in which there is 18% kerosene. k/(k+p) = .18 and the rest is p lol p/(k+p) = .82 k/.18 = (k+p) = p/.82 Eight litres are drawn off and then the vessel is filled with petrol. new amount: k+p-8+Lp that might be better

OpenStudy (amistre64):

so the new ratios are: k/(k+p-8+Lp) = .15 p/(k+p-8+Lp) = .85

OpenStudy (amistre64):

we can substitute one of the variables, p or k, and possible (k+p)

OpenStudy (amistre64):

k/.18 = (k+p) = p/.82 k+p = p/.82 k = 9/41 p k = .15(k+p-8+Lp) p = .85(k+p-8+Lp) 9/41 p = .15(p/.82-8+Lp) p = .85(p/.82-8+Lp)

OpenStudy (mathmath333):

i still dont get